OpenStudy (brynndelyn):

What is the square root of -72

Mehek (mehek14):

\(\Large\sqrt{-72}\) like this?

OpenStudy (asnaseer):

what do you think is the answer?

OpenStudy (brynndelyn):

6 sqrt 2i

OpenStudy (brynndelyn):

@asnaseer

OpenStudy (asnaseer):

perfect! :)

OpenStudy (brynndelyn):

i apparently got it wrong

OpenStudy (asnaseer):

I assume you are saying: \(6i\sqrt{2}\) i.e. the "i" in not inside the squareroot symbol

OpenStudy (asnaseer):

i.e. the "i" IS not inside...

OpenStudy (brynndelyn):

oh no, i was saying \[6\sqrt{2i}\]

OpenStudy (asnaseer):

ah - thats where you went wrong then... think of it like this:\[\sqrt{-72}=\sqrt{72}\times\sqrt{-1}=6\sqrt{2}\times i=6i\sqrt{2}\]

OpenStudy (brynndelyn):

oh okay that makes much more since, thanks!

OpenStudy (asnaseer):

yw :)